A ray of light is traveling in a glass with an index of refraction of n = 1.50 and strikes a glass/air interface. If the angle of incidence is 35°, Calculate the angle of refraction?


1
Expert's answer
2020-10-13T09:35:57-0400

Solution

In optics refractive index of the medium is given by

"n=\frac{sin \space i}{\sin r}"

Angle of incidence i=35°

Angle of refraction=r

Refractive index n=1.50

"\sin r=\frac{\sin 35°}{1.50}=0.38"

Angle of refraction

"r=\arcsin (0.38) =22.33°"